Find all the solutions to the equation 5x2 + 3x -1 = 7.
I put in 1 for x and got 5(1)^2 + 3(1) - 1 actually equal to 7, but it says find all solutions and I believe there's another solution.

Answer by ankor@dixie-net.com(22740)

5x^2 + 3x -1 = 7
5x^2 + 3x - 1 - 7 = 0
5x^2 + 3x - 8 = 0
Factors to
(5x+8)(x-1) = 0
Two solutions
x = 1
and
5x = -8
x = -8/5
x = -1.6
:
:
You can check the last solution in the original equation also.
5(-1.6^2) + 3(-1.6) - 1 = 7